# Boundary transfer matrices arising from quantum symmetric pairs

### v1

@@ 6.4. Spectral K-matrices
For any pseudo-involution $\phi:{\mathfrak{g}}\to{\mathfrak{g}}$, the automorphism $\phi_{q}$ descends to an automorphism of $U_{q}(\mathscr{L})$ and the subalgebra $U_{q}(\mathfrak{k})$ embeds into $U_{q}(\mathscr{L})$ as a right coideal subalgebra. We now follow [AV22b, Sec. 4] and observe that for the particular grading shift $\Sigma_{z}$ considered in Lemma 6.3.1, there exists a cylindrical structure $(\psi,J,K)$ on the affine QSP $(U_{q}({\mathfrak{g}}),U_{q}(\mathfrak{k}))$ such that
@@
Following the discussion in [AV22b, Secs. 5 and 6], we note that this intertwining condition is a consistent finite linear system with a 1-dimensional solution space if $V\in{\mathcal{C}}$ is irreducible as a $U_{q}({\mathfrak{g}}^{\prime})$-module. One obtains that $K_{V}(z)$ is proportional to a matrix $K^{\sf trig}_{V}(z)$ depending rationally on $z$, called trigonometric K-matrix. Both $K_{V}(z)$ and $K^{\sf trig}_{V}(z)$ satisfy the generalized spectral reflection equation
@@ 6.5. K-matrices for Kirillov-Reshetikhin modules
Recall that a Kirillov-Reshetikhin (KR) module is an irreducible object in ${\mathcal{C}}$ with trivial Drinfeld polynomials except for one, say associated to node $i\in\mathring{I}$, where the roots are given by a geometric progression in ${\mathbb{F}}$ with ratio $q_{i}$ (see e.g., [CP94, Her06] for more details). Suppose that $V,W\in{\mathcal{C}}$ are KR modules and that $\phi$ is restrictable: the subdiagram $\mathring{I}$ of finite type is $\tau$-stable and contains $X$. Then, by [AV22b, Thm. 7.8.1], in (6.19) $\psi$ can be replaced by a diagram automorphism.
The study of solutions of spectral reflection equations in various special cases, see e.g., [RV18, KOW22], suggests the following conjecture.
@@
Symmetrizable invertible solutions of (6.19) in tensor products of KR modules with $\psi$ equal to a diagram automorphism are rescaled actions of the grading-shifted universal K-matrix associated to a QSP with generalized parameter constraints as described in Section 5.4.
@@
Conversely, by [RV18] there are non-symmetrizable invertible solutions of (6.19), some of which cannot be obtained by taking limits of symmetrizable solutions. The general description of the corresponding coideal subalgebra $B\subset U_{q}({\mathfrak{g}})$ is an open question. It appears that only the triangular q-Onsager algebra, which embeds into $U_{q}(\widehat{\mathfrak{sl}}_{2})$, has been explicitly described, see [BB13]. $\triangledown$
@@ 6.6. Spectral tensor and dual K-matrices
in $\operatorname{End}({\mathsf{f}}_{\phi}\boxtimes{\mathsf{f}}_{{\mathcal{C}}}\boxtimes{\mathsf{f}}_{{\mathcal{C}}})((y,z/y))$, see [AV24, Sec. 6.4]. We use the following invertible element of $\operatorname{End}({\mathsf{f}}_{{\mathcal{C}}})((z))$ as a parameter-dependent gauge transformation:
@@
Crucially, note the spectral parameter shift in terms of $p=q^{-h_{\tau}^{\vee}}$. Now we define
@@
$({\operatorname{id}}\otimes\Delta_{z/y})(\overline{\mathbb{K}}(y))=(\widetilde{R}(yz)^{\overline{\psi}(y)})_{23}\cdot\overline{\mathbb{K}}(z)_{13}\cdot R(yz)^{\overline{\psi}(y)}_{23}\cdot\overline{\mathbb{K}}(y)_{12},$ (6.24)
@@
$=(\widetilde{K}(y)\otimes 1)\cdot\widetilde{R}(yz)^{\psi}\cdot(1\otimes\widetilde{K}(z))\cdot J.$ (6.29)
The identity (6.26) is easily seen to be equivalent to
@@
which is equivalent to Lemma 6.3.1. ∎
@@ 6.7. Boundary transfer matrices
Let $[{\mathcal{C}}]$ be the Grothendieck ring of finite-dimensional (type ${\bf 1}$) $U_{q}({\mathfrak{g}}^{\prime})$-modules.
@@
Let $(U_{q}({\mathfrak{g}}),U_{q}(\mathfrak{k}))$ be a QSP of affine type. There is a ring homomorphism $\mathscr{T}(z):[{\mathcal{C}}]\to\operatorname{End}({\mathsf{f}}_{\phi})\otimes{\mathbb{F}}((z))$ given by $\mathscr{T}^{(V)}(z)\coloneqq\mathop{\sf Tr}_{2}{\overline{{\mathbb{K}}}}_{\bullet,V}(z)$ for $V\in{\mathcal{C}}$. In particular, for all $V,W\in{\mathcal{C}}$, $[\mathscr{T}^{(V)}(y),\mathscr{T}^{(W)}(z)]=0$ in $\operatorname{End}({\mathsf{f}}_{\phi})\otimes{\mathbb{F}}((y,z))$.
@@
    In contrast with the case of finite type, $\mathscr{T}(z)$ is typically highly nontrivial as we show below in the case of the augmented q-Onsager algebra. $\triangledown$
6.8. Connection with quantum integrability
To see that Sklyanin’s formalism of commuting transfer matrices is recovered as a special case, as before we assume that $\phi$ is restrictable and the modules in question are KR modules, so that the twist automorphism $\psi$ appearing in (6.19) can be replaced by a diagram automorphism. If in addition $\phi$ satisfies the property that the longest element of $\mathring{W}=W(\mathring{{\mathfrak{g}}})$ sends $\alpha_{i}$ to $-\alpha_{\tau(i)}$ for all $i\in\mathring{I}$, then by [AV22b, Thm. 7.8.1] the untwisted spectral reflection equation holds, as originally considered in [Che84, Skl88]:
@@ 6.8. Connection with quantum integrability
For the spectral dual K-matrix defined by (6.22), we obtain (either from (6.31) and the group-like nature of $\mathring{D}$, or from (6.29) and the intertwining property of $R(z)$) the dual reflection equation, see [Skl88, (13)], [^16] This is up to conventions: a left-right swap needs to be made, and strictly speaking one obtains a mild generalization, see e.g., [Vla15, (2.13)]. which is interpreted as a statement of factorizable scattering at another (in this case, right) boundary
@@
$\qquad=\widetilde{K}_{V}(y)_{1}\cdot\widetilde{R}_{VW}(yz)\cdot\widetilde{K}_{W}(z)_{2}\cdot R_{W,V}(z/y)^{-1}_{21},$
Next, let $V_{1},\ldots,V_{N}\in{\mathcal{C}}$ be arbitrary and let $W$ be a KR module. By [AV24, Lem. 6.3.1], $V_{1}\otimes\cdots\otimes V_{N}$ is an object in $\mathcal{W}_{\phi}$. Therefore $\mathscr{T}^{(W)}(z)\in\operatorname{End}({\mathsf{f}}_{\phi})\otimes{\mathbb{F}}((z))$ acts on $V_{1}\otimes\cdots\otimes V_{N}$, yielding the linear map
@@
$E_{1},F_{0}\mapsto\begin{pmatrix}0&1\\ 0&0\end{pmatrix},\qquad F_{1},E_{0}\mapsto\begin{pmatrix}0&0\\ 1&0\end{pmatrix},\qquad k_{1},k_{0}^{-1}\mapsto\begin{pmatrix}q&0\\ 0&q^{-1}\end{pmatrix},$ (6.37)
which is a KR module satisfying $V^{\psi}=V$ where $\psi$ is a gauge transformation of $\phi_{q}$ by means of a Cartan correction. In this case, we have $\mathring{D}_{V}=\operatorname{Id}_{V}$. By solving the $\psi$-twisted $U_{q}(\mathfrak{k})$-intertwining condition, we readily obtain a well-known trigonometric K-matrix:
@@
Setting $z=1$ in $\frac{{\sf d}}{{\sf d}z}\mathscr{T}^{(V)}_{V^{\otimes N}}(z)$, one obtains the Hamiltonian for the XXZ spin chain with diagonal boundary conditions (with the two sets of QSP parameters related to each other). This is a highly nontrivial operator, making clear that $\mathscr{T}^{(V)}(z)$ is not a scalar. $\triangledown$
6.9. Representation-theoretic applications
We discuss some open questions surrounding the boundary transfer matrix, which would provide connections with various representation-theoretic constructions.
@@ 6.9. Representation-theoretic applications
    It is natural to interpret the image of $\mathscr{T}$ in $U_{q}(\mathfrak{k})\otimes{{\mathbb{C}}}(\negthinspace(z)\negthinspace)$ as a boundary analogue of a Bethe subalgebra. For example, in the case of (an extension of) the q-Onsager algebra, the expansion coefficients have been computed in [BK05, BB13, Lem23].
@@
    Fusion formulas for the boundary transfer matrices $\mathscr{T}^{(W)}_{V}(z)$ (TT relations), have been considered in special cases, originally by L. Mezincescu and R. Nepomechie in [MN92]. Using universal transfer matrices with auxiliary space equal to a spin-$j$ module of $U_{q}(\widehat{\mathfrak{sl}}_{2})$, this has been accomplished by P. Baseilhac, A. Gainutdinov and G. Lemarthe in [LBG23, Lem23] for the alternating central extension of the q-Onsager algebra $\mathcal{A}_{q}$, subject to conjectural identification of an explicitly constructed fused K-operator and the action of a tentative universal tensor K-matrix for $\mathcal{A}_{q}$. It would be nice to connect the universal element $\overline{\mathbb{K}}$ discussed here to the desired universal tensor K-matrix for $\mathcal{A}_{q}$, and approach boundary TT relations for other types follow from the universal transfer matrices discussed in this paper.
@@
    To upgrade the map $\mathscr{T}$ to a q-character map along the lines of [FR99, FM01] it is necessary to define a Harish-Chandra map. It is to be expected that it requires a Drinfeld-type loop presentation of $U_{q}(\mathfrak{k})$, which was recently obtained for various quasi-split types in the works [LW21, Zh22, LWZ23a, LWZ23b].
@@
    We obtained a ring homomorphism: $[{\mathcal{C}}]\to U_{q}(\mathfrak{k})\otimes{{\mathbb{C}}}(\negthinspace(z)\negthinspace)$ by letting the second leg of $\overline{\mathbb{K}}$ act on a quantum group module. One can instead evaluate the first leg of $\overline{\mathbb{K}}$, yielding a group homomorphism $[{\operatorname{Mod}}_{\sf fd}(U_{q}(\mathfrak{k}))]\to U_{q}({\mathfrak{g}}^{\prime})\otimes{{\mathbb{C}}}(\negthinspace(z)\negthinspace)$. After composition with a QSP analogue of a Harish-Chandra map $U_{q}(\mathfrak{k})\to U_{q}(\mathring{{\mathfrak{h}}}\otimes{\mathbb{C}}[t,t^{-1}])$, we expect this should reproduce the construction in [Prz23, LP24].
@@
    The quantum group $U_{q}({\mathfrak{g}})$ has two important categories of representations: ${{\mathcal{O}}^{\sf int}}$ and ${\mathcal{C}}$. They are related by intertwiners satisfying quantum KZ equations of type A, described in [FR92] in terms of its universal R-matrix $R$. We expect that, for a suitable choice of dual K-matrix, the tensor K-matrix $\overline{\mathbb{K}}$ of an affine QSP can be used to define analogous intertwiners connecting certain infinite-dimensional $U_{q}(\mathfrak{k})$-modules with objects in ${\mathcal{C}}$.
@@
    Finally, as mentioned in the introduction to [AV24], an overarching goal that connects with other points mentioned here, is to formulate a universal theory of Baxter relations for quantum integrable systems with reflecting boundaries. Given that $\overline{\mathbb{K}}$ lies in a completion of $U_{q}(\mathfrak{k})\otimes U_{q}({\mathfrak{b}}^{+})$ (at least in any quasi-split type, for a suitable dual K-matrix $\widetilde{K}$), it would be natural to follow the representation-theoretic approach of [HJ12, FH15].

### v2

@@ 6.4. Spectral K-matrices
For any pseudo-involution $\phi:{\mathfrak{g}}\to{\mathfrak{g}}$, the automorphism $\phi_{q}$ descends to an automorphism of $U_{q}(\mathscr{L})$ and the subalgebra $U_{q}(\mathfrak{k})$ embeds into $U_{q}(\mathscr{L})$ as a right coideal subalgebra. We now follow [AV25a, Sec. 4] and observe that for the particular grading shift $\Sigma_{z}$ considered in Lemma 6.3.1, there exists a cylindrical structure $(\psi,J,K)$ on the affine QSP $(U_{q}({\mathfrak{g}}),U_{q}(\mathfrak{k}))$ such that
@@
Following the discussion in [AV25a, Secs. 5 and 6], we note that this intertwining condition is a consistent finite linear system with a 1-dimensional solution space if $V\in{\mathcal{C}}$ is irreducible as a $U_{q}({\mathfrak{g}}^{\prime})$-module. One obtains that $K_{V}(z)$ is proportional to a matrix $K^{\sf trig}_{V}(z)$ depending rationally on $z$, called trigonometric K-matrix. Both $K_{V}(z)$ and $K^{\sf trig}_{V}(z)$ satisfy the generalized spectral reflection equation
@@ 6.5. K-matrices for Kirillov-Reshetikhin modules
Recall that a Kirillov-Reshetikhin (KR) module is an irreducible object in $\mathcal{C}$} with trivial Drinfeld polynomials except for one, say associated to node $i\in\mathring{I}$, where the roots are given by a geometric progression in $\mathbb{F}$} with ratio $q_{i}$ (see, e.g., [CP94, Her06] for more details). Suppose that $V,W\in{\mathcal{C}}$ are KR modules and that $\phi$ is $\tau$-restrictable, i.e., the subdiagram $\mathring{I}$ of finite type is $\tau$-stable (or, equivalently, $\tau$ fizes the affine node). Then, up to a gauge transformation, $K_{V}(z)$ and $K_{W}(z)$ satisfy a generalized spectral reflection equation replace (6.18) where $\psi$ is replaced by a diagram automorphism, see [AV25a, Thm. 7.8.1].
The study of solutions of spectral reflection equations in various special cases, see, e.g., [RV18, KOW22], suggests the following conjecture.
@@
Consider the equation (6.18) with $\psi$ replaced by a diagram automorphism. Symmetrizable invertible solutions of this equation in tensor products of KR modules are rescaled actions of the grading-shifted universal K-matrix associated to a QSP of affine type with generalized parameter constraints as described in Section 5.4.
@@
Conversely, by [RV18] there are non-symmetrizable invertible solutions of (6.18), some of which cannot be obtained by taking limits of symmetrizable solutions. The general description of the corresponding coideal subalgebras $B\subset U_{q}({\mathfrak{g}})$ is an open question. It appears that only the triangular q-Onsager algebra, which embeds into $U_{q}(\widehat{\mathfrak{sl}}_{2})$, has been explicitly described, see [BB13]. $\triangledown$
@@ 6.6. Spectral tensor and dual K-matrices
in $\operatorname{End}({\mathsf{f}}_{\phi}\boxtimes{\mathsf{f}}_{{\mathcal{C}}}\boxtimes{\mathsf{f}}_{{\mathcal{C}}})((y,z/y))$, see [AV25b, Sec. 6.4]. We use the following invertible element of $\operatorname{End}({\mathsf{f}}_{{\mathcal{C}}})((z))$ as a parameter-dependent gauge transformation:
@@
Crucially, note the spectral parameter shift in terms of $p=q^{h_{\phi}^{\vee}}$. Now we define
@@
$({\operatorname{id}}\otimes\Delta_{z/y})(\overline{\mathbb{K}}(y))=\widetilde{R}(yz)^{\overline{\psi}(y)}_{23}\cdot\overline{\mathbb{K}}(z)_{13}\cdot R(yz)^{\overline{\psi}(y)}_{23}\cdot\overline{\mathbb{K}}(y)_{12},$ (6.23)
@@
$=(\widetilde{K}(y)\otimes 1)\cdot\widetilde{R}(yz)^{\psi}\cdot(1\otimes\widetilde{K}(z))\cdot J,$ (6.28)
where we relied on the $p$-shift appearing in the definition of $\widetilde{K}(z)$. The identity (6.25) is easily seen to be equivalent to
@@
which in turn is equivalent to Lemma 6.3.1. ∎
@@ 6.7. Boundary transfer matrices
Let $\mathcal{C}$}] be the Grothendieck ring of finite-dimensional (type $\bf 1$}) $U_{q}({\mathfrak{g}}^{\prime})$-modules.
@@
Let $(U_{q}({\mathfrak{g}}),U_{q}(\mathfrak{k}))$ be a QSP of affine type. There is a ring homomorphism $\mathscr{T}(z):[{\mathcal{C}}]\to\operatorname{End}({\mathsf{f}}_{\phi})\otimes{\mathbb{F}}((z))$ given by $\mathscr{T}^{(V)}(z)\coloneqq\mathop{\sf Tr}_{2}\overline{\mathbb{K}}_{\bullet,V}(z)$ for $V\in{\mathcal{C}}$. In particular, for all $V,W\in{\mathcal{C}}$, $\mathscr{T}^{(V)}(y),\mathscr{T}^{(W)}(z)]=0$ in $\operatorname{End}({\mathsf{f}}_{\phi})\otimes{\mathbb{F}}((y,z))$.
@@
    Even if $K=K^{\prime}$, in contrast with the case of finite type, $\mathscr{T}(z)$ is typically highly nontrivial as we show below in the case of the augmented q-Onsager algebra. $\triangledown$
6.8. Sklyanin’s construction
We explain how Sklyanin’s formalism of commuting transfer matrices can be recovered as a special case. Suppose $V,W\in{\mathcal{C}}$ satisfy $V^{\psi}=V$ and $W^{\psi}=W$ (in [AV25a, Thm. 7.8.1], it is shown that if $\phi$ is $\tau$-restrictable, the longest element of the Weyl group of $\mathring{{\mathfrak{g}}}$ sends $\alpha_{i}$ to $\alpha_{\tau(i)}$ for all $i\in\mathring{I}$ and $V$ is a KR-module, then $V^{\psi}=V$). Then the untwisted spectral reflection equation holds, as originally considered in [Che84, Skl88]:
@@ 6.8. Sklyanin’s construction
For the spectral dual K-matrix defined by (6.21), we obtain (either from (6.30) and the group-like nature of $\mathring{D}$, or from (6.28) and the intertwining property of $R(z)$) the following dual reflection equation:
@@
$\qquad=\widetilde{K}_{V}(y)_{1}\cdot\widetilde{R}_{VW}(yz)\cdot\widetilde{K}_{W}(z)_{2}\cdot R_{W,V}(z/y)^{-1}_{21}.$
This equation appeared in [Skl88, (13)], [^16] This is up to conventions: an overall left-right swap needs to be made, and strictly speaking one obtains a mild generalization, see, e.g., [Vla15, (2.13)]. with the interpretation of factorizable scattering at a second (in our case, right) boundary.
Next, let $V_{1},\ldots,V_{N},W\in{\mathcal{C}}$ be arbitrary. In quantum integrability, the tensor product $V_{1}\otimes\cdots\otimes V_{N}$ is interpreted as the space of states of a quantum spin chain and $W$ is called auxiliary space. By [AV25b, Lem. 6.3.1], $V_{1}\otimes\cdots\otimes V_{N}$ is an object in $\mathcal{W}_{\phi}$. Therefore $\mathscr{T}^{(W)}(z)\in\operatorname{End}({\mathsf{f}}_{\phi})\otimes{\mathbb{F}}((z))$ acts on $V_{1}\otimes\cdots\otimes V_{N}$, yielding the linear map
@@
$E_{1},F_{0}\mapsto\begin{pmatrix}0&1\\ 0&0\end{pmatrix},\qquad F_{1},E_{0}\mapsto\begin{pmatrix}0&0\\ 1&0\end{pmatrix},\qquad k_{1},k_{0}^{-1}\mapsto\begin{pmatrix}q&0\\ 0&q^{-1}\end{pmatrix}$. (6.36)
We can now explicitly check that $V^{\psi}=V$ where $\psi$ is a gauge transformation of $\phi_{q}$ by a Cartan correction. In this case, we have $\mathring{D}_{V}=\operatorname{Id}_{V}$. By solving the $\psi$-twisted $U_{q}(\mathfrak{k})$-intertwining condition, we readily obtain a well-known trigonometric K-matrix:
@@
Setting $z=1$ in $\frac{{\sf d}}{{\sf d}z}\mathscr{T}^{(V)}_{V^{\otimes N}}(z)$, one obtains the Hamiltonian for the XXZ spin chain with diagonal boundary conditions (with the two sets of QSP parameters related to each other). This is a highly nontrivial operator, making clear that $\mathscr{T}^{(V)}(z)$ is not a scalar matrix. $\triangledown$
6.9. Connections
We discuss some, generally speculative, connections of the boundary transfer matrix with various representation-theoretic constructions.
@@ 6.9. Connections
    It is natural to interpret the image of $\mathscr{T}(z)$ in $U_{q}(\mathfrak{k})\otimes{{\mathbb{C}}}(\negthinspace(z)\negthinspace)$ as a boundary analogue of a Bethe subalgebra. For example, in the case of (an extension of) the q-Onsager algebra, the expansion coefficients have been computed in [BK05, BB13, Lem23].
@@
    Fusion formulas for the boundary transfer matrices $\mathscr{T}^{(W)}_{V}(z)$ (TT relations), have been considered in special cases, originally by L. Mezincescu and R. Nepomechie in [MN92]. Using universal transfer matrices with auxiliary space equal to a spin-$j$ module of $U_{q}(\widehat{\mathfrak{sl}}_{2})$, this has been accomplished by G. Lemarthe in [Lem23, Thm. 4.2.5], also see [LBG25b], for the alternating central extension $\mathcal{A}_{q}$ of the q-Onsager algebra, subject to conjectural identification of an explicitly constructed fused K-operator and the action of a tentative universal tensor K-matrix for $\mathcal{A}_{q}$. It would be nice to connect the universal element $\overline{\mathbb{K}}$ to the desired universal tensor K-matrix for $\mathcal{A}_{q}$, and approach boundary TT relations for other types from the universal transfer matrices discussed in this paper.
@@
    To upgrade the map $\mathscr{T}(z)$ to a q-character map along the lines of [FR99, FM01] it is necessary to define a QSP analogue of the loop Harish-Chandra map. This requires a Drinfeld-type loop presentation of $U_{q}(\mathfrak{k})$, which was recently obtained for various quasi-split types in the works [LW21, Zh22, LWZ23, LWZ24].
@@
    We obtained a ring homomorphism: ${\mathcal{C}}]\to U_{q}(\mathfrak{k})\otimes{{\mathbb{C}}}(\negthinspace(z)\negthinspace)$ by letting the second leg of $\overline{\mathbb{K}}$ act on a quantum group module. One can instead evaluate the first leg of $\overline{\mathbb{K}}$, yielding a group homomorphism ${\operatorname{Mod}}_{\sf fd}(U_{q}(\mathfrak{k}))]\to U_{q}({\mathfrak{g}}^{\prime})\otimes{{\mathbb{C}}}(\negthinspace(z)\negthinspace)$. When $X$ does not contain the affine node, we expect it to generalize the boundary q-character map defined in [Prz25, LP24] in quasi-split types, after composition with the loop Harish-Chandra map.
@@
    In affine type, the quantum group $U_{q}({\mathfrak{g}})$ has two important categories of representations: ${\mathcal{O}}^{\sf int}$} and $\mathcal{C}$}. They are related by intertwiners satisfying quantum KZ equations of type A, described in [FR92] in terms of its universal R-matrix $R$. We expect that, for a suitable choice of dual K-matrix, the tensor K-matrix $\overline{\mathbb{K}}$ of an affine QSP can be used to define analogous intertwiners connecting certain infinite-dimensional $U_{q}(\mathfrak{k})$-modules with objects in $\mathcal{C}$}.
@@
    Finally, as mentioned in the introduction to [AV25b], an overarching goal that connects with other points mentioned here, is to formulate a universal description of spectra for quantum integrable systems with reflecting boundaries. In the case of systems on a circle (periodic boundary conditions), the central idea is that the spectra are determined by so-called Baxter relations, which should be viewed as relations in the Grothendieck ring of a suitable category of $U_{q}({\mathfrak{b}}^{+})$-modules, see [HJ12, FH15]. It is natural to suspect a similar theory can be developed for systems with reflecting boundary conditions.
    We can make the following convenient observation, which means that $\mathscr{T}^{(W)}(z)$ can be defined for suitable $U_{q}({\mathfrak{b}}^{+})$-modules for certain affine QSPs. Suppose that the dual K-matrix $\widetilde{K}$ appearing in $\overline{\mathbb{K}}$ is itself obtained, see Remark 6.7.2, from the universal basic K-matrix $K^{\prime}$ associated to an affine quantum symmetric pair. Further, suppose that the affine quantum symmetric pair defining the basic K-matrix $K$ is quasi-split. If both $K$ and $K^{\prime}$ are part of standard cylindrical structures, then indeed $\overline{\mathbb{K}}$ lies in a completion of $U_{q}(\mathfrak{k})\otimes U_{q}({\mathfrak{b}}^{+})$.